# Control Strategies for Pursuit-Evasion Under Occlusion Using Visibility and Safety Barrier Functions

### v1

@@ I INTRODUCTION
While early works on CBFs primarily focused on continuous systems subject to differentiable CBF constraints, recent advances [22, 23, 18] have studied the regularity of CBF-based controllers and have expanded their applicability to non-smooth dynamics and non-smooth constraints. These method have been used in [24] to enable safe multi-robot navigation with polytope-shaped robots. In our case, the visibility CBF is Lipschitz continuous but not differentiable everywhere. It is also time-varying due to the pursuer’s and evader’s motion. Hence, it is necessary to use a non-smooth time-varying formulation of the visibility CBF constraint involving its generalized gradient.
Maintaining visibility is a requirement in various target tracking problems. A visibility-aware planner proposed in [25] generates sensor trajectories with maximal visibility to a moving target. A vision-based controller for a quadrotor landing on a ground vehicle, while preventing camera occlusions, was developed in [26]. Leader-follower formations for maintaining visibility and safety were developed using dipolar vector fields in a known environment in [27]. Similarly, in the presence of FoV constraints, dipolar vector fields have been used in [28] to formulate a model predictive control scheme for differential-drive robot navigation, maintaining visibility of a static landmark. Closely related to our work, Gao et al. [29] formulate a probabilistic notion of visibility under occlusion, which is maintained using an extended Kalman filter [30]. The authors develop a real-time non-myopic trajectory planner for visibility-aware and safe target tracking in the presence of uncertainty. This work is complementary to ours in that it focuses on high-level non-myopic visibility planning using nonlinear optimization, while we focus on low-level myopic visibility control to track a planned reference using convex programming. The contributions of our work are summarized as follows.
@@ III-A Control Barrier Functions
$\mathcal{S}(t)=\{\mathbf{x}\in\mathbb{R}^{n}|\;h(t,\mathbf{x})\geq 0\},$ (2)
@@ III-C Planning and Control with Visibility & Safety Constraints
To increase the efficiency of SST, we used the Motion Planning Transformer (MPT) [37] to suggest sampling regions. MPT uses a transformer architecture that maps patches of an occupancy grid map to a sequence of regions that may contain feasible trajectories to the goal. SST’s sampling is biased to the sequence of regions predicted by MPT to focus the tree expansion towards areas that may observe the evader.
@@ IV-A CARLA Simulations
Ablation Planner only Controller only Our Method
@@
We also carried out an ablation study of the planner and controller components in our method with different pursuer and evader initialization. We compared: (a) directly using the planned reference trajectory to control the pursuer (Planner only); (b) using only the low-level controller with visibility CBF constraint (Controller only); and (c) our method. The results are reported in Table I. Planner only struggled without real-time safety constraints, leading to collisions. Controller only performed better than Planner only, maintaining visibility and avoiding collisions. Our method shows the best performance in visibility and safety across all metrics, maintaining the evader within the pursuer’s FoV for 98% of the trajectory while avoiding collisions with obstacles. This demonstrates the efficacy of our approach in providing more reliable and safer autonomous navigation in simulated environments that closely mimic real-world conditions.
@@ IV-B Real World Experiments
In the real experiments, the pursuer was a differential-drive Jackal robot, equipped with an OS-1 Ouster 3D LiDAR and D455 RealSense depth camera. We used two types of evaders: a human and a TurtleBot robot. An AprilTag was affixed to the evader to aid in camera detections. We imposed a limited, triangular FoV $\mathcal{F}_{0}$ which spans $30^{\circ}$ with range $2$ m to demonstrate our method’s ability to deal with limited sensing. We ran the Controller Only method onboard the Jackal, consisting of sensing, mapping, evader detection, and control. We omitted the planner for computational reasons, allowing us to run the controller at 50 Hz, and from simulation results we see that the Controller Only method achieves fast, safe tracking. The pursuer control space $\mathcal{U}$ was set to $[0,0.5]$ m/s $\times$ $[-0.5,0.5]$ rad/s. We conducted all experiments in a lab environment and used common objects as obstacles.
@@ V CONCLUSION
This paper introduces a convex programming formulation for control synthesis with visibility and safety constraints. We show that both constraints can be formulated using non-smooth but Lipschitz continuous CBFs. Coupling our controller with a kinodynamic planner that generates non-myopic trajectories to observe a mobile evader, we demonstrate autonomous robot pursuit in 3D simulations and real-world environments using onboard sensing under occlusion. Future work will focus on supplementing our method with an exploration strategy in the event that the pursuer does not begin with a detection of the evader, and with an estimator that maintains multiple hypotheses for evader motion.

### v2

@@ I INTRODUCTION
While early works on CBFs primarily focused on continuous systems subject to differentiable CBF constraints, recent advances [22, 23, 18] have studied the regularity of CBF-based controllers and have expanded their applicability to non-smooth dynamics and non-smooth constraints. These methods have been used in [24] to enable safe multi-robot navigation with polytope-shaped robots. In our case, the visibility CBF is Lipschitz continuous but not differentiable everywhere. It is also time-varying due to the pursuer’s and evader’s motion. Hence, it is necessary to use a non-smooth time-varying formulation of the visibility CBF constraint involving its generalized gradient.
Maintaining visibility is a requirement in various target tracking problems. A visibility-aware planner proposed in [25] generates sensor trajectories with maximal visibility to a moving target. A vision-based controller for a quadrotor landing on a ground vehicle, while preventing camera occlusions, was developed in [26]. Leader-follower formations for maintaining visibility and safety were developed using dipolar vector fields in a known environment in [27]. Similarly, in the presence of FoV constraints, dipolar vector fields have been used in [28] to formulate a model predictive control scheme for differential-drive robot navigation, maintaining visibility of a static landmark. Closely related to our work, Gao et al. [29] formulate a probabilistic notion of visibility under occlusion, which is maintained using an extended Kalman filter [30]. The authors develop a real-time non-myopic trajectory planner for visibility-aware and safe target tracking under uncertainty. This work is complementary to ours in that it focuses on high-level non-myopic visibility planning using nonlinear optimization and linear SDF approximations, while we focus on low-level myopic visibility control to track a planned reference using convex programming. The contributions of our work are summarized,
@@ III-A Control Barrier Functions
$\mathcal{S}(t)$ =$=\{\mathbf{x}\in\mathbb{R}^{n}|\;h(t,\mathbf{x})\geq 0\},$ (2)
@@ III-C Planning and Control with Visibility & Safety Constraints
To increase the efficiency of SST, we use the pre-trained Motion Planning Transformer (MPT) [37] to suggest sampling regions. MPT uses a transformer architecture that maps patches of an occupancy grid map to a sequence of regions that may contain feasible trajectories to the goal. SST’s sampling is biased to the sequence of regions predicted by MPT to focus the tree expansion towards areas that may observe the evader.
@@ IV-A CARLA Simulations
Ablation Planner only Controller only Full system
@@
We also carried out an ablation study of the planner and controller components in our method with different pursuer and evader initialization. We compared: (a) directly using the planned reference trajectory to control the pursuer (Planner only); (b) using only the controller with visibility and safety CBF constraints (Controller only); and (c) Asynchronous planner+controller (Full system). The results are reported in Table I. Planner only struggled without real-time safety constraints, leading to collisions. Controller only performed better than Planner only, maintaining visibility and avoiding collisions. Full system shows the best performance in visibility and safety across all metrics, maintaining the evader within the pursuer’s FoV for 98% of the trajectory while avoiding collisions with obstacles. This shows our method’s effectiveness for reliable, safe navigation in realistic simulated environments.
@@ IV-B Real World Experiments
In the real experiments, the pursuer was a differential-drive Jackal robot, equipped with an OS-1 Ouster 3D LiDAR and D455 RealSense depth camera. We used a human-controlled Turtlebot, affixed with an AprilTag as an evader. We imposed a limited, triangular FoV $\mathcal{F}_{0}$ which spans $30^{\circ}$ with range $2$ m to demonstrate our method’s ability to deal with limited sensing. We ran the Controller Only method at 50 Hz onboard the Jackal, consisting of sensing, mapping, evader detection, and control. Due to the complex environment and high computational demands (about ten times slower for MPT inference), we excluded the planner on Jackal robot, which struggles with incomplete and noisy maps to provide a reliable path. The pursuer control space $\mathcal{U}$ was set to $[0,0.5]$ m/s $\times$ $[-0.5,0.5]$ rad/s. We conducted all experiments in a lab environment and used common objects as obstacles.
@@ V CONCLUSION
This paper introduces a convex programming formulation for control synthesis with visibility and safety constraints. We show that both constraints can be formulated using non-smooth but Lipschitz continuous CBFs. Coupling our controller with a kinodynamic planner that generates non-myopic trajectories to observe a mobile evader, we demonstrate autonomous robot pursuit in 3D simulations and real-world environments using onboard sensing under occlusion. Future work will focus on supplementing our method with an exploration strategy if the pursuer does not begin with a detection of the evader, and with a robust estimator that maintains multiple hypotheses for evader motion.